answersLogoWhite

0

🎒

Probability

The probability of a certain event is a number expressing the likelihood that a specific event will occur, expressed as the ratio of the number of actual occurrences to the number of possible occurrences. In mathematics, it is a measure of how often an event will happen and is the basis of statistics.

14,643 Questions

You roll a dice once what is the probability that you roll an odd number?

there is six sides on a dice and three of them is odd

3/6=1/2=50%

it's a 50% chance to roll an odd number

A women is going to have 6 children What is the probability that all 6 will be boys?

if we assume that the probability for a girl being born is the same as a boy being born: (1/2)^6 = 0.015625 = 1.5625%

What compares the probability of an event occurring to the probability that it will not occur?

If the probability of an event occurring is p, then 1-p represents the probability of the same event not occurring. The value of p must lie between 0 and 1.

How do you calculate Z and T scores?

z=(x-mean)/(standard deviation of population distribution/square root of sample size)

T-score is for when you don't have pop. standard deviation and must use sample s.d. as a substitute.

t=(x-mean)/(standard deviation of sampling distribution/square root of sample size)

What does standard precautions mean?

Washing hands after patient contact; using gloves when touching blood, body fluids, secretions, excretions, and contaminated items;.

What is the probability of drawing the jack of hearts then the queen of hearts?

It depends on what exactly the conditions are. If you mean from a 52 card deck, drawing one card right after the other, then the chance that they will be those two cards in that exact order is 1/52 * 1/51. This is pretty much a percent of a percent. .00038, or about 4 percent of a percent.

What is the probability of getting all tails if a coin is tossed 3 times?

There is a one in two chance of tails every time, rewritten as 1/2

The probability of tail 3 times in a row is 1/2 x 1/2 x 1/2= 1/8.

So the answer is one in eight, or 12.5%

What is the probability of getting 3 heads in 10 tosses of a fair coin?

130 instances of 3 heads out of 1024 total possible outcomes=130/1024=0.126953125

What is the probability of getting 6 heads in 10 tosses of a fair coin?

(10C6)[P(H)^4][P(T)^6]

= (210)[(1/2)^4][(1/2)^6]

= (210)[(1/2)^10]

= 0.205 In case you are not familiar with the notation, (10C6) means 10 choose 6. It is the number of way to pick 6 things from 10 when the order does not matter.

We calculate this as 10C6 = 10!/[(10 - 6)!6!] = 10!/(4!6!) = (10 x 9 x 8 x 7x 6!)/(4 x 3 x 2 x 1 x 6!) simplify

= 10 x 3 x 7 = 210

If two cards are simultaneously drawn from a standard deck of cards what is the probability that they are not both the same suit?

The first card can be anything. That means that there are 51 cards left and 39 of them are not the same suit as the first one, therefore P(not same suit) = 39/51 = 13/17.

If a coin is flipped 7 times what is the probability of landing on all tails?

each time you flip the coin, probability to end on either side is 50% (or 0.5) (we disregard landing on the side).

So, to land on the same side 7 times, it is: 0.5^7